# A Two-Stage Stochastic Model for Road-Rail Intermodal Freight Transportation Under Demand and Capacity Uncertainty

### v1

@@ 4.3 Deterministic equivalent of stochastic optimization model
$+\lambda\text{CVaR}_{\alpha}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)+\sum_{\omega\in\omega}\rho\eta_{\omega}$ (20)
@@
$+\lambda\text{CVaR}+\sum_{t\in T}\rho\eta_{\omega}$ (22)
@@
Where $\xi_{\omega}$ denote the excess cost by which the cost exceeds the $\theta$ in each scenario and $c_{e\omega}$ represent the total transportation and penalty costs for every scenario. (23) ensures that the excess cost is calculated correctly based on the difference between the total cost in each scenario and the VaR. Constraint (24) ensures that excess is non-negative. While (24) defines CVaR as the VaR ($\theta$) plus the expected excess cost over all scenarios, scaled by $\frac{1}{1-\alpha}$. CVaR represents the expected value of the excess cost beyond the $\theta$, providing a measure of the average loss in the worst $(1-\alpha)\%$ of cases.
@@ 5 Case Study and Computational Results
This section applies the proposed two-stage stochastic optimization model to an intermodal transportation problem in the U.S. The analysis utilizes publicly available datasets, including FAF and NTAD, which provide detailed information on rail connections, intermodal terminals, and road links to hubs from supply locations such as warehouses and manufacturing facilities. To address data gaps, missing supply-hub connections were supplemented by manually retrieving data from Google Maps. Figures 2(a) and 2(b) provide a visual overview of the spatial layout of the U.S. rail network, highlighting connectivity between supply points, intermodal hubs, and key corridors. This foundational data supports the development of a network model, detailed subsequently.
@@
The computational experiments were executed on a machine equipped with a 12th Gen Intel(R) Core(TM) i7-1265U processor, 16 GB RAM, and a 64-bit operating system. Each instance was solved independently using the Gurobi Optimizer 11.0.0, ensuring efficient handling of the model’s complexity. The implementation was performed in Python 3.11.4. The results of these experiments, which offer critical insights into the model’s behavior under real-world constraints, are detailed in the subsequent subsections.
@@ 5.1 Policy Questions
    Balancing Cost, Risk, and Decision-Making Under Uncertainty: How does the model navigate trade-offs between cost efficiency and risk mitigation, particularly under varying values of the risk-aversion parameter and confidence level? How do these risk-based decisions affect cost structures, system resilience, and operational priorities across different stakeholder preferences? Additionally, how does risk aversion influence the stability and predictability of cost outcomes across multiple uncertainty scenarios?
@@
    Impact of Capacity Constraints on Cost and Demand Satisfaction: How do train capacity limitations affect total system costs and unmet demand, and at what point do additional capacity increases yield diminishing returns?
@@
    Impact of Risk Preferences on Cost Stability and Scenario Variability: How does increasing $\lambda$ affect the variability of cost outcomes across different demand scenarios? At what point do additional risk-mitigation efforts yield diminishing returns in stabilizing total costs?
By addressing these questions, the case study evaluates the model’s ability to optimize intermodal logistics under uncertainty while balancing cost minimization, risk resilience, and sustainability objectives. The subsequent sections provide detailed analyses and insights into these aspects, demonstrating the model’s practical applicability and contributions to both research and real-world industry practices.
@@ 5.2 Case Study Setup
To ensure uniformity across the rail network, FedEx warehouses were used as the supply locations. This choice was made to standardize the analysis and maintain consistency in evaluating logistics across different regions. FedEx warehouses, known for their strategic distribution capabilities, provide a practical and representative basis for modeling supply points in intermodal transportation. Furthermore, some of these warehouses serviced multiple intermodal hubs, reflecting the flexibility and extensive reach of distribution networks in real-world logistics.
To further enhance the accuracy and applicability of the model, some parameter values were adopted from peer-reviewed works. For example, the planning time horizon was similar to the 60 hours (or 2.5 days) used in Ke and Verma, (2021). Additionally, Assadipour et al., (2016) estimated that approximately one hour is required to load, unload, or transfer an intermodal container, a value incorporated into the model to reflect realistic handling times at intermodal hubs. These parameter choices ensure the model aligns with observed industry practices and remains grounded in empirical evidence, balancing cost efficiency and operational feasibility.
@@ 5.2.1 Network Dynamics
Building on the case study setup, this section details the operational framework of the intermodal transportation network, focusing on train and road transportation dynamics. These elements are critical for understanding how the proposed model optimizes logistics under real-world constraints.
  • Train Operations: The train operations modeled in this study are based on fixed schedules with defined origin-destination pairs and intermediate stops. Below, we describe the operations of one representative train (Train A), which connects major industrial hubs and serves as a vital corridor for intermodal freight movement:
    • Route Description: Train A begins its journey in Buffalo, NY, and terminates in Detroit, MI (Delray), traversing key economic regions. Intermediate stops include Greencastle, PA; Columbus, OH; Cincinnati, OH; St. Louis, MO; and Decatur, IL.
    • Operational Characteristics: The total route covers approximately 1,642 miles, with each intermediate stop allocated six hours of dwell time to facilitate loading, unloading, and container transfers. These dwell times are integral to aligning train schedules with supply and road network connections.
    • Additional Routes: Other intermodal routes in the network are detailed in the appendix (Tables A.1, A.2, and A.3), reflecting diverse O-D pairs and operational constraints across the modeled system. These routes contribute to evaluating the scalability of the model.
  • Road Transportation: The road network forms a critical connection to the intermodal system. Each supply location is linked to its nearest intermodal hub based on a fixed distance metric, ensuring efficiency in routing. While dynamic routing is not considered in this study, this assumption reflects industry practices where companies rely on pre-determined routes for recurring freight movements. This fixed routing approach facilitates accurate cost and time estimates while maintaining alignment with real-world logistical operations.
@@ 5.3 Computational Experiment
To ensure clarity, several abbreviations are used throughout this section. OBJ denotes the stochastic model’s objective function, incorporating both first-stage and second-stage costs. E(TC) represents the expected total cost across all scenarios, while ASC refers to the average stochastic cost, capturing the mean realized costs across experiments. CVaR (Conditional Value-at-Risk) quantifies the expected cost of extreme scenarios beyond a specified threshold, determined by $\alpha$.
The risk-aversion parameter $\lambda$ controls the trade-off between expected cost minimization and risk aversion. It is varied from 0 to 1 in increments of 0.1 (i.e., $\lambda=0,0.1,0.2,\ldots,1$), allowing us to explore the full spectrum of decision-making priorities, from purely cost-focused strategies ($\lambda=0$) to highly risk-averse approaches ($\lambda=1$). Similarly, the risk confidence level $\alpha$ is tested at values of 0.25, 0.5, 0.75, and 1. These values provide insights into the model’s response to varying levels of risk tolerance, from focusing on extreme losses in a quarter of scenarios ($\alpha=0.25$) to considering worst-case outcomes ($\alpha=1$).
@@
5.3.1 Impact of Risk-Aversion Parameter and Confidence Level on Cost and CVaR
The interplay between $\lambda$ and $\alpha$ provides valuable insights into how the model balances total cost minimization and risk mitigation. Understanding how these parameters influences cost and risk is crucial in decision-making under uncertainty. The parameter $\lambda$ dictates the weight placed on the CVaR, effectively controlling the trade-off between minimizing average costs and reducing exposure to extreme cost scenarios. Meanwhile, $\alpha$ determines the quantile of the cost distribution used for CVaR calculations, thereby capturing the confidence level for considering risk reflecting the level of conservatism in risk management.
@@ 5.3.1 Impact of Risk-Aversion Parameter and Confidence Level on Cost and CVaR
Table 2: Results for different $\alpha$ and $\lambda$ values (scaled to $10^{3}$).
@@
Table 2 and Figure 3(a) illustrate the impact of increasing $\lambda$ on the objective value. As expected, higher $\lambda$ values result in a steady increase in the objective function across all $\alpha$ levels. This outcome is intuitive since a greater emphasis on extreme scenarios leads to higher overall costs, as the model prioritizes robust solutions that minimize worst-case losses. However, the rate of increase in cost varies with $\alpha$, with higher confidence levels (e.g., $\alpha=0.90$) showing a steeper rise compared to lower ones (e.g., $\alpha=0.25$). This suggests that risk-averse settings amplify the effect of risk mitigation efforts, making them increasingly expensive as confidence levels increases.
Conversely, Table 2 and Figure 3(b) show the diminishing CVaR values as $\lambda$ increases. The reduction in CVaR demonstrates the model’s ability to shift resources toward solutions that mitigate extreme cost realizations. However, beyond a certain point ($\lambda\approx 0.5$), additional increases in $\lambda$ yield only marginal improvements in risk reduction, indicating diminishing returns in CVaR minimization. Notably, this effect is more pronounced for higher $\alpha$ values, reinforcing the idea that extreme risk management strategies become increasingly costly and less effective beyond a specific risk-aversion threshold.
@@
  • Growing Cost Disparity Across $\alpha$ Levels: The widening gap in objective values for different $\alpha$ settings at higher $\lambda$ levels indicates that high-risk-averse settings amplify the sensitivity of the model to changes in confidence levels.
@@
Implications for Decision-Makers
These findings offer valuable insights for strategic decision-making:
@@
  • Optimal Risk Balancing Strategy: The presence of diminishing returns in CVaR reduction suggests that a moderate risk-aversion setting ($\lambda\approx 0.5$) may offer the best trade-off between cost efficiency and worst-case scenario protection.
In summary, the results emphasize the importance of balancing cost efficiency with resilience against extreme cost events. The widening gap in objective values across $\alpha$ levels suggests that models incorporating dynamic risk-preference adjustments can better align with diverse operational realities. Future work could explore how additional constraints, such as capacity limitations, supply fluctuations, and intermodal transfer restrictions, interact with risk-based cost optimizations.
@@ 5.3.2 Analysis of Cost Components
To gain a holistic understanding of the cost structure within the optimization framework, we analyze the aggregated cost components across all scenarios. The primary cost components include: Supply Cost, which represents the expenses associated with sourcing decisions; Transportation Cost, costs incurred in moving freight; Unmet Demand Penalty, which quantifies costs associated with unmet demand; and Emissions Penalty, reflecting the additional cost of exceeding predefined emissions limits. This analysis addresses Policy Question 1 by demonstrating how the model balances cost, risk, and decision-making under uncertainty, optimizing resource allocation to minimize costs while maintaining service levels and sustainability compliance. It also provides insights into Policy Question 2, showing how sustainability constraints affect cost distributions, and Policy Question 6, which examines the impact of capacity limitations on total costs and unmet demand.
Cost Drivers and Trade-offs in Optimization
The cost distribution, illustrated in Figure 5, is shaped by the model’s objective function and key operational constraints. The interplay between transportation, supply, unmet demand, and emissions penalties highlights the trade-offs involved in optimizing intermodal logistics.
  • Transportation Costs (58.1%) dominate total expenditures, reflecting the high logistical costs of intermodal freight movements. While the model prioritizes cost minimization, transportation remains the primary driver due to network topology and train capacity limits. To mitigate expenses, the model optimally assigns containers to the nearest supply locations, minimizing per-container transportation costs.
  • Supply Costs (34.8%) are influenced by sourcing decisions. Since supply nodes have finite capacities, the model balances cost-efficient procurement with constraints related to hub accessibility and demand fulfillment. Selecting lower-cost suppliers may increase emissions penalties or create mismatches between supply locations and demand hubs.
  • Unmet Demand Penalties (4.7%) arise from capacity constraints. The model strategically reduces unmet demand to minimize associated penalties, ensuring that available transportation capacity is used efficiently. However, limiting freight movement in low-demand scenarios may lower transportation costs while increasing unmet demand penalties.
  • Emissions Penalties (2.4%) remain relatively small, as the model integrates emissions constraints as a soft penalty rather than an absolute limit. While the current configuration allows minor exceedances, enforcing stricter sustainability targets would shift costs toward transportation and supply expenses.
The results underscore the interdependencies between cost components—minimizing one often leads to trade-offs
@@
Implications for Decision-Makers
The cost breakdown provides actionable insights for stakeholders:
  • Optimizing transportation operations is crucial, as it represents the largest cost component. Efficient route selection, train scheduling, and capacity planning can significantly reduce expenses by minimizing unnecessary freight movements and optimizing modal shifts. Additionally, reducing inefficiencies in transshipment operations and leveraging real-time data for dynamic scheduling can further cut costs. Stakeholders must also consider congestion effects at hubs, as delays and rerouting can indirectly increase transportation expenses.
  • Sustainability constraints have a minimal cost impact, reinforcing the model’s ability to balance emissions reduction with economic feasibility. The low emissions penalties suggest that the model effectively guides stakeholders toward environmentally responsible logistics without imposing excessive costs. Optimizing routes and sourcing locations inherently reduces emissions while maintaining cost efficiency, making the model a valuable tool for companies aiming to meet carbon reduction targets without sacrificing profitability.
  • Capacity constraints affect total costs, especially under demand uncertainty. While increasing capacity can reduce unmet demand penalties, the model demonstrates that dynamic allocation and smarter load balancing can achieve similar benefits without major infrastructure investments. Rather than expanding physical capacity, decision-makers should focus on adaptive planning, demand forecasting, and flexible routing to minimize service disruptions and cost escalations.
Broader Research Implications
This analysis highlights the importance of an integrated optimization approach, where transportation efficiency, sustainability, and capacity planning are addressed collectively rather than as independent factors. The results demonstrate the model’s effectiveness in supporting logistics providers, policymakers, and industry stakeholders in developing balanced, cost-efficient, and sustainable intermodal transportation strategies. Future research could explore the incorporation of dynamic pricing mechanisms to mitigate unmet demand penalties and evaluate alternative routing strategies to further optimize transportation costs. These findings provide a foundation for strategic policy development in intermodal freight planning, ensuring both economic viability and environmental sustainability in logistics operations.

### v2

@@ 4.3 Deterministic equivalent of stochastic optimization model
$+\lambda\text{CVaR}_{\alpha}\left(\sum_{i\in I}\sum_{j\in J}\sum_{n\in N}\sum_{t\in T}C_{ij}x_{ijn\omega t}+\pi_{\omega}\sum_{n\in N}U_{n\omega}\right)+\rho\sum_{\omega\in\Omega}p_{\omega}\eta_{\omega}$ (21)
@@
$+\lambda\text{CVaR}+\rho\sum_{\omega\in\Omega}p_{\omega}\eta_{\omega}$ (23)
@@
Where $\xi_{\omega}$ denote the excess cost by which the cost exceeds the $\theta$ and $c_{e\omega}$ represent the total transportation and penalty costs for every scenario. (24) ensures that the excess cost is calculated correctly based on the difference between the total cost in each scenario and the VaR. Constraint (25) ensures that excess is non-negative. While (25) defines CVaR as the VaR ($\theta$) plus the expected excess cost over all scenarios, scaled by $\frac{1}{1-\alpha}$. CVaR represents the expected value of the excess cost beyond the $\theta$, providing a measure of the average loss in the worst $(1-\alpha)\%$ of cases.
@@ 5 Case Study and Computational Results
This section applies the proposed two-stage stochastic optimization model to an intermodal transportation problem in the U.S. The analysis utilizes publicly available datasets, including FAF U.S. Department of Transportation, 2024a and NTAD U.S. Department of Transportation, 2024b , which provide detailed information on rail connections, intermodal terminals, and road links to hubs from supply locations such as warehouses and manufacturing facilities. To address data gaps, missing supply-hub connections were supplemented by manually retrieving data from Google Maps. Figures 2(a) and 2(b) provide a visual overview of the spatial layout of the U.S. rail network, highlighting connectivity between supply points, intermodal hubs, and key corridors. This foundational data supports the development of a network model, detailed subsequently.
@@
The computational experiments were executed on a machine equipped with a 12th Gen Intel(R) Core(TM) i7-1265U processor, 16 GB RAM, and a 64-bit operating system. Each instance was solved independently using the Gurobi Optimizer 11.0.0, ensuring efficient handling of the model’s complexity. The implementation was performed in Python 3.11.4. The results of these experiments are detailed in the subsequent subsections.
@@ 5.1 Policy Questions
    Balancing Cost, Risk, and Decision-Making Under Uncertainty: How does the model manage trade-offs between cost efficiency and risk mitigation under varying levels of risk aversion and confidence? What are the implications for system resilience, cost structure, and stakeholder priorities? How does increasing risk aversion affect the stability and predictability of outcomes across scenarios?
@@
    Impact of Capacity Constraints on Cost and Demand Satisfaction: How do train capacity affect total system costs and unmet demand, and at what point do additional capacity increases yield diminishing returns?
@@
    Impact of Risk Preferences on Cost Stability and Scenario Variability: How does increasing $\lambda$ affect cost outcomes across demand scenarios? Do additional risk-mitigation efforts yield diminishing returns?
By addressing these questions, the case study evaluates the model’s ability to optimize intermodal logistics under uncertainty while balancing cost minimization, risk resilience, and sustainability objectives. The subsequent sections provide detailed analyses into these aspects, demonstrating the model’s practical applicability and contributions.
@@ 5.2 Case Study Setup
To ensure uniformity across the rail network, FedEx warehouses were used as supply locations. This choice was made to standardize the analysis and maintain consistency in evaluating logistics across different regions. FedEx warehouses, known for their strategic distribution capabilities, provide a practical and representative basis for modeling supply points in intermodal transportation. Furthermore, some of these warehouses serviced multiple intermodal hubs, reflecting the flexibility and extensive reach of distribution networks in real-world logistics.
Parameter values are drawn from peer-reviewed studies and public sources to ensure realism and relevance. $C_{ij}$ and $\pi_{\omega}$ follow intermodal cost structures in (Goodarzi et al.,, 2024), while $O_{i}$ is set to \$250/container, similar to estimates in (FHWA,, 2025). The emissions cap $\epsilon$ adopts the approach in (Heinold and Meisel,, 2020), representing the share of emissions a shipper is willing to tolerate, and is varied in our analysis. $E_{t}$ uses modal emission factors provided by EPA guidance (Agency,, 2024). The emissions penalty $\rho$ is set to \$20 per metric ton of CO2, reflecting typical social cost benchmarks. Additionally, to enhance realism, the planning horizon (60 hours) is aligned with (Ke and Verma,, 2021), and a 1-hour per-container handling time at hubs is based on estimates from (Assadipour et al.,, 2016).
@@ 5.2.1 Network Dynamics
This study models an intermodal freight transport network comprising five scheduled train routes spanning major U.S. economic corridors. Each route connects strategic hubs via fixed schedules, facilitating container transfers between rail and road. Train operations are characterized by defined origin–destination pairs, intermediate stops, and fixed dwell times of approximately six hours per stop to allow for loading, unloading, and intermodal handling.
As an illustrative example, Train A travels from Buffalo, NY, to Detroit, MI (Delray), covering  2,105 miles. It stops at Greencastle, Columbus, Cincinnati, St. Louis, and Decatur. Each segment has calibrated travel times based on actual rail distances, with intermediate dwell periods to align with realistic freight operations. The routing structure reflects real-world transit logistics and timing constraints. For instance, the Buffalo–Greencastle segment spans 280 miles and 6.5 hours, while Decatur–Detroit covers 310 miles over 7 hours.
Each hub is linked to multiple nearby facilities (e.g., warehouses or plants), with distances typically under 100 miles. These are selected based on proximity and relevance to freight demand. For example, FedEx Warehouse 13 is 68 miles from Buffalo, while facilities near St. Louis range from 11 to 29 miles away. These first-mile connections are assumed to use fixed road routes, aligning with typical industry practices. Road segments are modeled using average speeds (55 mph) and costs (\$3/mile), enabling precise estimation of total delivery time and cost from origin to hub.
Other routes (Trains B–E) span cities such as Seattle, Dallas, Jacksonville, and New Orleans, incorporating 4–5 intermediate hubs each, which can be recreated. While Train A is detailed for clarity, the others follow similar structures, and their inclusion demonstrates the model’s scalability across network configurations.
@@ 5.3 Computational Experiment
To ensure clarity, several abbreviations are used throughout this section. OBJ denotes the stochastic model’s objective function, incorporating both first-stage and second-stage costs. E(TC) represents the expected total cost across all scenarios, while ASC refers to the average stochastic cost, capturing the mean realized costs across experiments. CVaR quantifies the expected cost of extreme scenarios beyond a specified threshold, determined by $\alpha$.
The parameter $\lambda$ controls trade-off between expected cost minimization and risk aversion. It is varied from 0 to 1 in increments of 0.1 (i.e., $\lambda=0,0.1,0.2,\ldots,1$), allowing us to explore the full spectrum of decision-making priorities, from cost-focused strategies ($\lambda=0$) to highly risk-averse approaches ($\lambda=1$). Similarly, the risk confidence level $\alpha$ is tested at values of 0.25, 0.5, 0.75, and 1. These values provide insights into the model’s response to varying levels of risk tolerance, from focusing on extreme losses in a quarter of scenarios ($\alpha=0.25$) to worst-case outcomes ($\alpha=1$).
@@
5.3.1 Impact of Risk-Aversion Parameter ($\lambda$) and Confidence Level ($\alpha$) on Cost and CVaR
The interplay between $\lambda$ and $\alpha$ provides valuable insights into how the model balances total cost and risk. Understanding the effect of these parameters is crucial in decision-making under uncertainty. The parameter $\lambda$ dictates the weight placed on the CVaR, controlling the trade-off between minimizing average costs and reducing exposure to extreme cost scenarios. Meanwhile, $\alpha$ determines the quantile of the cost distribution used for CVaR calculations, thereby capturing the confidence level for considering risk reflecting the level of conservatism in risk management.
@@ 5.3.1 Impact of Risk-Aversion Parameter ($\lambda$) and Confidence Level ($\alpha$) on Cost and CVaR
Table 2: Results for different $\alpha$ and $\lambda$ values (costs in \$ $\times$ 103)
@@
Table 2 and Figure 3(a) illustrate the impact of increasing $\lambda$ on the objective. As expected, higher $\lambda$ result in a steady increase in the objective across all $\alpha$ levels. This outcome is intuitive since a greater emphasis on extreme scenarios leads to higher overall costs, as the model prioritizes robust solutions that minimize worst-case losses. However, the rate of increase in cost varies with $\alpha$, with higher confidence levels (e.g., $\alpha=0.90$) showing a steeper rise compared to lower ones (e.g., $\alpha=0.25$). This suggests that risk-averse settings amplify the effect of risk mitigation efforts, making them increasingly expensive as confidence levels increases.
Conversely, Table 2 and Figure 3(b) show the diminishing CVaR values as $\lambda$ increases. The reduction in CVaR demonstrates the model’s ability to shift resources toward solutions that mitigate extreme cost realizations. However, beyond a certain point ($\lambda\approx 0.5$), additional increasing $\lambda$ yield only marginal improvements in risk reduction, indicating diminishing returns in CVaR minimization. Notably, this effect is more pronounced for higher $\alpha$ values, reinforcing the idea that extreme risk management strategies become more costly and less effective beyond a specific threshold.
@@
  • Growing Cost Disparity Across $\alpha$ Levels: The widening gap in objective values for different $\alpha$ values at higher $\lambda$ levels indicates that high-risk-averse settings amplify sensitivity of the model to changes in confidence levels.
@@
paragraphImplications for Decision-Makers These findings offer valuable insights for strategic decision-making:
@@
  • Optimal Risk Balancing Strategy: The presence of diminishing returns in CVaR reduction suggests that a moderate risk-aversion ($\lambda\approx 0.5$) may offer the best trade-off between cost and worst-case scenario protection.
In summary, the results emphasize the importance of balancing cost efficiency with resilience against extreme cost events. The widening gap in objective values across $\alpha$ levels suggests that models incorporating dynamic risk-preference adjustments can better align with diverse operational realities.
5.3.2 Insights from VSS and EVPI Across Problem Scales
To assess the benefit of modeling uncertainty, we compute the Value of the Stochastic Solution (VSS) and the Expected Value of Perfect Information (EVPI) across varying problem sizes, defined by combinations of train counts and scenario sets (Table 3). The table summarizes several key performance metrics: the EEV column reports the expected cost of applying the deterministic plan under uncertainty; SS gives the cost of the optimal stochastic solution; and WS reflects the cost assuming perfect foresight (i.e., full scenario realization). The VSS quantifies the benefit of adopting a stochastic model over a deterministic one, while EVPI measures the hypothetical gain from perfect information. The normalized columns VSS (%) scale these gains relative to the stochastic solution cost. Lastly, CVaRSS denotes the Conditional Value-at-Risk, capturing expected cost in the worst-case tail under the stochastic policy.
The results show that VSS values are consistently positive, confirming the advantage of stochastic optimization over deterministic planning. For example, with 4 trains and 6 scenarios, the VSS indicates a 3.8% cost reduction. This advantage tends to increase with more scenarios, particularly in smaller configurations where deterministic policies exhibit greater vulnerability to uncertainty.
EVPI values, however, are substantially higher than VSS in all settings, e.g., \$52.71k in the (4,8) instance, underscoring the large theoretical value of perfect foresight. The persistent gap between EVPI and VSS reflects the limitations of nonanticipative models, which must commit to first-stage decisions before uncertainty is resolved.
Table 3: Impact of Stochastic Optimization and Perfect Information Across Problem Sizes (Costs in \$ $\times$ 103)
Trains Scen EEV $S$ WS VSS EVPI VSS (%) CVaRSS
3 3 44.90 43.69 14.69 1.21 29.00 2.8 31.25
3 6 63.16 61.50 19.09 1.66 42.41 2.7 48.10
3 8 51.11 50.31 13.26 0.80 37.05 1.6 33.62
4 3 50.52 49.27 19.17 1.25 30.10 2.5 32.35
4 6 71.86 69.20 22.82 2.66 46.38 3.8 50.30
4 8 76.16 74.66 21.95 1.50 52.71 2.0 50.54
7 3 89.09 88.82 33.03 0.27 55.80 0.3 56.35
7 6 126.82 121.90 40.84 4.91 81.06 4.0 100.60
7 8 110.22 105.97 35.49 4.25 70.48 4.0 82.40
CVaR values remain proportionally reasonable relative to the SS cost, suggesting that the model effectively controls tail risk without overly inflating expected expenditures. Key takeaways: (i) stochastic optimization yields measurable gains over deterministic plans; (ii) perfect information offers significantly greater cost-saving potential; and (iii) increasing system flexibility (via more trains) tends to reduce the marginal value of stochastic modeling.
Computationally, solving the model across these scales remains tractable for small to moderate instances. The two-stage structure with binary first-stage and integer second-stage decisions, along with CVaR risk terms and inter-scenario coupling, renders the problem NP-hard due to their combinatorial structure and the need to jointly evaluate all scenarios. To assess practical solvability, for the smallest case (3 trains, 3 scenarios), the solve time was under 10 seconds. As the problem scaled to 7 trains and 8 scenarios, solve times increased significantly, with some instances taking several hours to reach optimality. These results demonstrate that the model is tractable for moderate problem sizes, but larger real-world applications may benefit from tailored decomposition or heuristic approaches.
@@ 5.3.3 Analysis of Cost Components
To better understand the cost structure within our optimization framework, we analyze the aggregated cost components across all scenarios: Supply Cost, Transportation Cost, Unmet Demand Penalty, and Emissions Penalty. This breakdown addresses Policy Questions 1, 2, and 6 by illustrating how the model balances economic efficiency, sustainability, and operational constraints under uncertainty.
Cost Drivers and Trade-offs
Figure 5 shows that transportation costs dominate (58.1%), driven by the logistical demands of container movement across the network. Despite the model’s cost-minimization objective, train capacity constraints and network topology elevate these costs. Supply costs (34.8%) reflect sourcing decisions, constrained by location capacities and their proximity to demand nodes.
Unmet demand penalties (4.7%) stem from capacity shortfalls, especially under peak demand scenarios. The model prioritizes fulfilling demand where feasible but accepts limited shortfalls to avoid disproportionate cost escalation. Emissions penalties (2.4%), while minor, signal the model’s ability to respect sustainability constraints without inflating total costs, though stricter limits could shift this balance.
@@
Insights for Decision-Makers
The cost breakdown offers several strategic insights:
  • Transportation efficiency is critical, as it represents the largest share of total costs. Optimizing routes, schedules, and hub operations especially under real-time constraints can yield substantial savings.
  • Sustainability goals can be met at low cost. The low emissions penalties indicate that the model supports carbon-conscious logistics with minimal financial trade-off, aligning environmental and economic objectives.
  • Capacity constraints drive unmet demand, but the model shows that smart allocation and flexible sourcing often outperform infrastructure expansion. Adaptive planning and accurate forecasting are key to managing costs under uncertainty.
Broader Implications
This analysis reinforces the value of an integrated planning framework that accounts for operational, economic, and environmental factors. By internalizing emissions and unmet demand into the cost function, the model enables balanced, data-driven decision-making for sustainable intermodal logistics. Future extensions could incorporate dynamic pricing or more granular emissions tracking to further refine trade-off management.